OpenStudy (anonymous):

If 30.0 L of oxygen are cooled from 200ºC to 1°C at constant pressure, what is the new volume of oxygen? (Can you pls show the steps)

OpenStudy (abmon98):

Volume and temperature, if we heat the gas particles, they move more rapidly gas expands!. this means they have a directly proportional relationship at constant number of moles and pressure. use Charles Law V1/V2=T1/T2

OpenStudy (abmon98):

Directly proportional relationships means that when volume is decreases temperature is cooled and when the temperature increases volume expands. 30.00/V2=200/1 V2*200=30.00*1 V2=30.00/200.00